In southern California, a growing number of individuals pursuing teaching credentials are choosing paid internships over traditional student teaching programs. A group of eight candidates for



three local teaching positions consisted of five who had enrolled in paid internships and three who



enrolled in traditional student teaching programs. All eight candidates appear to be equally qualified, so three are randomly selected to fill the open positions. Let Y be the number of internship



trained candidates who are hired.



(a) Does Y have a binomial or hypergeometric distribution? Why?



(b) Find the probability that two or more internship trained candidates are hired



(c) What are the mean and standard deviation of Y ?.




1
Expert's answer
2022-03-28T10:49:51-0400

"a:\\hypergeometric\,\,because\,\,the\,\,candidates\,\,are\,\,chosen\,\,without\,\,replacement\\parameters\,\,N=8,K=5,n=3\\b:\\P\left( Y\geqslant 2 \right) =P\left( Y=2 \right) +P\left( Y=3 \right) =\frac{C_{5}^{2}C_{3}^{1}}{C_{8}^{3}}+\frac{C_{5}^{3}C_{3}^{0}}{C_{8}^{3}}=0.714286\\c:\\EY=n\frac{K}{N}=3\cdot \frac{5}{8}=1.875\\
\sigma Y=\sqrt{\frac{nK\left( N-K \right) \left( N-n \right)}{N^2\left( N-1 \right)}}=\sqrt{\frac{3\cdot 5\left( 8-5 \right) \left( 8-3 \right)}{8^2\left( 8-1 \right)}}=0.708683"